OpenStudy (anonymous):

-9(y-15)-20=-3y-(y-5) have to solve for y

OpenStudy (anonymous):

-9y+135-20=-4y+5 5y=110 y=22

OpenStudy (anonymous):

-9y+135-20=-3y-y+5 -9y+115=-4y+5 -9y+4y=5-115 -5y=-110 y=-110/-5 y=22